OpenStudy (anonymous):

After how many months will Rita pay two hundred dollars in simple annual interest on a credit card that charges fifteen percent on her four hundred dollars balance?

OpenStudy (anonymous):

@agent0smith @ganeshie8 @Luigi0210 @jdoe0001

OpenStudy (anonymous):

@eliassaab @uri @genius12

OpenStudy (anonymous):

@aprilann143

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@agent0smith @whpalmer4 @ganeshie8 @robtobey @eliassaab @Luigi0210 @Compassionate @Destinymasha @uri @surjithayer @kewlgeek555 HEEEELP!

OpenStudy (anonymous):

interest=p*r*t/100 \[200=\frac{ 400*15*t }{ 100 },t~ is~ \in~ years\]

OpenStudy (anonymous):

T is what?

OpenStudy (anonymous):

t is time.

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):

yeah t is time p is principal i is intrest r is rate

OpenStudy (anonymous):

So what do I do?

OpenStudy (anonymous):

i=p*r*t

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

\[t=\frac{ 200 }{60 }=\frac{ 10 }{ 3 }*12=40~ months.\]

OpenStudy (anonymous):

but 40 months is how many years?

OpenStudy (whpalmer4):

You could probably work this out in your head. 15% annual simple interest is 15%/12 = 1.25% simple interest each month, or $1.25/$100 balance * $400 balance = $5/mo. $200/($5/month) = 40 months.

OpenStudy (whpalmer4):

1 year = 12 months 1 year/12 months = 1 40 months * (1 year/12 months) = 40/12 years = (you can reduce the fraction or make a decimal)

OpenStudy (whpalmer4):

problem does ask for "how many months", I note...

OpenStudy (anonymous):

so once you find the time in years you apply the number to the formula and BAM!!! you will have your answer

OpenStudy (whpalmer4):

no, you work it in months because you have monthly interest.

OpenStudy (anonymous):

ok maybe I was wrong but I'm only in 6th grade so...

OpenStudy (whpalmer4):

Well, you could work it in years as well, since you have simple interest in this problem. If the problem dealt with compound interest, you couldn't. Anyhow, any questions about how to get the answer?